# Phonon effects, impact ionization and power conversion in Mott photovoltaic systems

### v1

@@ III.2 Observables
with $t_{\text{rel}}=t-t^{\prime}$ and $t_{\text{av}}=(t+t^{\prime})/2$ the relative and average times, respectively. The mean $j$  [^3] As discussed in Ref. [31], in a steady-state situation $j_{\tilde{z},\tilde{z}+1}$ should have the same value for every $\tilde{z}$. Nevertheless, due to the limited accuracy of our AMEA impurity solver [40], deviations occur. For this reason, we consider the mean $j$ of the photocurrent (averaged over the bonds), together with its standard deviation $\sigma_{j}$. of the steady-state photocurrent flowing from the left to the right lead [40, 36, 31] and its corresponding standard deviation $\sigma_{j}$ are
$\begin{split}j&\equiv\frac{1}{\text{L}+1}\sum_{\tilde{z}=0}^{\text{L}}j_{\tilde{z},\tilde{z}+1},\\ \sigma_{j}&\equiv\sqrt{\sum_{\tilde{z}=0}^{\text{L}}\frac{1}{\text{L}+1}(j_{\tilde{z},\tilde{z}+1}-j)^{2}},\end{split}$ (15)
where $j_{\tilde{z},\tilde{z}+1}$ is the photocurrent between any of the $(\text{L}+1)$ bonds of the entire Mott photovoltaic device [^4] The bonds take into account also the one from the left lead ($\tilde{z}=0$) to the first layer ($\tilde{z}=1$) and from the last layer ($\tilde{z}=\text{L}$) to the right lead ($\tilde{z}=\text{L}+1$)., given by:
$j_{\tilde{z},\tilde{z}+1}=t_{\tilde{z},\tilde{z}+1}\int_{-\infty}^{+\infty}\frac{\mathop{}\!\mathrm{d}\omega}{2\pi}\int\mathop{}\!\mathrm{d}\epsilon\int\mathop{}\!\mathrm{d}\overline{\epsilon}\rho\operatorname*{Re}(J_{00,\tilde{z},\tilde{z}+1}),$ (16)
where $\rho$  [^5] In Eq. 16 we omit the ($\omega,\epsilon,\varepsilon$) dependence to facilitate the reading. is the JDOS from Sec. II and the index $\tilde{z}$ runs over the leads’ surface layers and the correlated region, i.e. $\tilde{z}\in\left\{0,1,\dots,\text{L}\right\}$ with $\tilde{z}=0$ indicating the left and $\tilde{z}+1=\text{L}+1$ the right lead surface, the GF of which are given in Eq. (10).
We characterize the spectral properties of this system via the local electron density of states (DOS) $A_{z}(\omega)$ and occupation function $N_{z}(\omega)$  [^6] The occupation function $N_{z}(\omega)$ tells us how many electronic states per spin are occupied with energy $\omega$ in the layer $z$ and is proportional to $G^{<}_{\text{loc},00,zz}$., defined as
$A_{z}(\omega)=-\frac{1}{\pi}\operatorname*{Im}[G_{\text{loc},00,zz}^{\text{R}}(\omega)],$ (17)
@@
$N_{z}(\omega)=\frac{1}{4\pi}\left\{\operatorname*{Im}\left[G^{\text{K}}_{\text{loc},00,zz}(\omega)\right]-2\operatorname*{Im}\left[G^{\text{R}}_{\text{loc},00,zz}(\omega)\right]\right\},$ (18)
in which $G^{\text{R}}_{\text{loc},00,zz}(\omega)$ and $G^{\text{K}}_{\text{loc},00,zz}(\omega)$ represent respectively the time-averaged retarded and Keldysh local GF, introduced in Appendix A. Another important quantity we consider is the layer-dependent double occupation
$N_{\text{D},z}=\langle\hat{n}_{z,\uparrow}\hat{n}_{z,\downarrow}\rangle,$ (19)
@@
$\begin{split}\Delta N_{\text{D},z}&=\langle\left(\hat{n}_{z,\uparrow}-\langle\hat{n}_{z,\uparrow}\rangle\right)\left(\hat{n}_{z,\downarrow}-\langle\hat{n}_{z,\downarrow}\rangle\right)\rangle\\ &=N_{\text{D},z}-n_{z}^{2}\end{split}$ (20)
and a mean-field contribution $n_{z}^{2}$, with $n_{z}=\langle\hat{n}_{z,\uparrow}\rangle=\langle\hat{n}_{z,\downarrow}\rangle$ denoting the number of particles per spin per layer  [^7] In this work, we consider the central correlated system in a paramagnetic phase. As mentioned in Sec. I, antiferromagnetic correlations are important for the spreading of photoexcited carriers [38] and should be included in the model for a possible enhancement of the efficiency of a Mott photovoltaic system.. $n_{z}$ enters also into the total exceeding charge in the layer $z$
@@
A quantity related to the efficiency of the Mott photovoltaic device, which is discussed in Sec. IV, is the time-averaged absorbed power $\overline{P}_{\text{abs}}$
$\begin{split}&\overline{P}_{\text{abs}}=-E_{0}\int\mathop{}\!\mathrm{d}\epsilon\int\mathop{}\!\mathrm{d}\overline{\epsilon}\rho(\epsilon,\overline{\epsilon})\sum_{l=-\infty}^{\infty}\Big{(}J_{1-l}(\mathscr{A})\\ &+J_{-1-l}(\mathscr{A})\Big{)}\Big{[}\epsilon(1-\delta_{l,0})\delta_{l,2k}+\operatorname*{i}\overline{\epsilon}\delta_{l,2k+1}\Big{]}\\ &\sum_{z=1}^{\text{L}}\int_{-\infty}^{+\infty}\frac{\mathop{}\!\mathrm{d}\omega}{2\pi}G_{l,z}^{<}(\omega;\epsilon,\overline{\epsilon})\end{split}$ (22)
in which $k\in\mathbb{Z}$. We use in (LABEL:eq:avg_abs_pow) the so-called Wigner representation, introduced in Appendix D, for the lesser GF (see Sec. III.1).
IV Results
We consider the setup shown in Fig. 1, with the local DOS landscape in Fig. 2, consisting of $\text{L}=4$ correlated layers, a situation which has already been studied in Refs. [45, 24, 31]. The default values for the main parameters employed in this manuscript are given in Table 1  [^8] The values of the parameters in Table 1 are the same of Ref. [31] and are such that $\alpha\equiv t^{\ast}_{\parallel}E_{0}/2\Omega^{2}<0.5$, which justifies FDSA [25, 30, 31]., if not stated otherwise.
$U$ $E_{0}$ $\gamma_{\text{l}/\text{r}}$ $v$ $t_{\text{l}/\text{r}}$ $t_{\perp}$ $\Phi$ $1/\beta$ L
@@ IV Results
Table 1: Default values of the main parameters used in this manuscript: for simplicity we defined $v\equiv v_{\text{l}/\text{r}}$. As a reminder, the renormalized hopping in the correlated region is $t^{\ast}_{\parallel}=2$.
@@
As in Ref. [31], we choose the leads’ onsite energies $\varepsilon_{\text{l}/\text{r}}$ such that the left(right) lead DOS overlaps with the lower(upper) Hubbard band of the leftmost(righmost) layer with $\varepsilon_{\text{l}}=\varepsilon_{1}$ and $\varepsilon_{\text{r}}=-\varepsilon_{\text{l}}$, see the scheme in Fig. 2.
@@
Figure 2: Schematic representation of impact ionization processes occurring in the system under analysis. Electrons are injected into the central insulating region (in green) by the left reservoir (in blue) and drained out of it by the right reservoir. $\Phi$ is the potential drop across the central region and is a linear function of the coordinate $z$ of the layers. The bias voltage $\Delta V=\mu_{r}-\mu_{l}$ applied to the metallic leads sets the occupations within them and produces a current from the right to the left of the heterostructure (orange arrows). For values of the driving frequency $\Omega\gtrsim 2\Delta_{\text{g}}$, a photoexcited electron in the upper band can excite, in turn, a second electron across the gap from the lower band by transferring energy to it via Coulomb scattering. As outcome, two carriers per photon can now tunnel towards the right lead, resulting in a current from the left to the right (denoted by the red arrows), which flows against the potential energy established by the bias voltage.
@@
We vary the bias voltage $\Delta V=\mu_{r}-\mu_{l}$, so that for $\Delta V>0$ energy is harvested from the periodic driving when an electron current flows from the left to the right lead. We take, as convention, the current flowing from right to left as positive and the one from left to right as negative. The local DOS of the correlated layers has a gap $\Delta_{\text{g}}\approx 3$, while the bandwidth of the leads $W_{\text{b}}\approx 8.5$ [30, 31]. We quantify the effect of acoustic phonons on the electronic observables via the effective e-ph coupling strength $\lambda_{\textrm{eff}}=2g^{2}/W_{\text{b}}\omega_{ph}$, whose derivation is found in Appendix C  [^9] The values of $\lambda_{\textrm{eff}}$ in Sec. IV are obtained taking $W_{\text{b}}=8.5$, $g\in\left\{0.3,0.4,0.5\right\}$ and $\omega_{ph}\in\left\{0.05,0.1\right\}$, in line with Ref. [30, 64]..
@@
In this work we are interested in assessing how II is influenced by the dissipation by acoustic phonons and how it affects the efficiency of the Mott photovoltaic system. This process, in which an electron excited to the upper Hubbard band (UHB) by a photon with energy $\Omega$ excites a second electron across the gap via electron-electron scattering, may occur if the bandwidth of the UHB, which in our setup is roughly equal to the leads’ DOS, is at least twice the size of the gap $\Delta_{\text{g}}$, i.e. $W_{\text{b}}\geq 2\Delta_{\text{g}}$ [25, 30, 31].
IV.1 Effects of electron-phonon interaction
@@
Figure 3: (a) Absolute value of the time-averaged steady state current $|j|$ with error bars corresponding to $\sigma_{j}$ and (b)-(c) fluctuation of double occupancy $\Delta N_{\text{D},z}$ for $z=1$ and $z=2$, plotted as functions of the driving frequency $\Omega$, for different values of the bias voltage $\Delta V$ without and with e-ph interaction with coupling strength $\lambda_{\text{eff}}=0.752$. Default parameters are specified in Table 1. Due to the PhI symmetry (see Ref. [31]), $\Delta N_{\text{D},\text{L}+1-z}=\Delta N_{\text{D},z}$ so the curves for $z=3$ and $z=4$ in (b)-(c) are omitted. (Here $\phi=2$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
@@
Fig. 3 shows that the photocurrent as well as $\Delta N_{\text{D},z}$ for $\Delta V=2$ in presence of the e-ph interaction (green curves) display a slight enhancement for small to intermediate $\Omega$-values. This is due to the tiny redistribution of the spectral weight from the maxima peaks towards the edges of the Hubbard bands in Figs. 4, which reduces slightly the gap $\Delta_{g}$, with a larger occupation of the positive-energy states next to the bottom of the UHBs for all the layers. This effect is relatively small for a single layer but becomes appreciable when considering the whole heterostructure, and leads to a small increase of the photocurrent, as already observed in Ref. [30]. The magnitude of the peak in $j$ at $\Delta V=2$ remains instead comparable, because of the weaker DOS suppression around the Hubbard peaks [Fig. 4(b)] with respect to Ref. [30].
@@
We notice in Fig. 3(a) that the kink around $\Omega\approx 2\Delta_{g}$ signalizing II (see Ref. [31]) is not affected by the e-ph interaction. In this regard, also the fluctuation of double occupancy $\Delta N_{\text{D},z}$ [green curves in Fig. 3(b)-(c)] shows a change in slope for $\Omega\approx 2\Delta_{g}$, although less sharp due to the smearing by acoustic phonons.
@@
On the other hand, at $\Delta V=5$ the photocurrent experiences a reduction in magnitude because of the larger bias voltage which favours a current in the opposite direction  [^10] The current at $\Omega\gtrsim 17$ for $\Delta V=5$ becomes negative (not displayed in Fig. 3(a)), because for such $\Omega$ the carriers’ flow goes in the direction of the bias voltage.. The presence of acoustic phonons suppresses further $j$ because the Hartree shift, discussed in Sec. IV.1.1, moves the Hubbard bands away from each other, and the overlap between the layers’ DOS is reduced, as shown in Fig. 5(a).
@@
For small $\Omega$ the available states in the layers close to the right lead are mostly filled (see $N_{z}(\omega)$ for $z=3,4$ in Fig. 5(a)). Therefore, they suppress the photocurrent and contribute to a substantial increase of $\Delta N_{\text{D},z}$ (coral curve in Fig. 3(b)-(c))  [^11] The residual value of $\Delta N_{\text{D},z}$ at small and large $\Omega$ (red and coral curves in Fig. 3(b)-(c)) is due to the large separation between the Hubbard bands as a consequence of the Hartree shift..
@@
Figure 4: (a) Spectral functions $A_{z}(\omega)$ (solid lines) and occupation functions $N_{z}(\omega)$ (shaded areas and dashdotted lines) for the different layers at $\Omega=8$, $\Delta V=2$ and $\lambda_{\textrm{eff}}=0.752$. (b) Zoom of electron DOS and occupation function for the layer $z=1$. Vertical dashed and dashdotted lines in (a) highlight the separation between the center of the bands of the layers $z=1$ and $z=4$, without (blue line) and with e-ph interaction (dark green line). Vertical black lines in (b) highlight the difference in the peak positions of the bands without and with e-ph interaction. Default parameters are specified in Table 1. (Here $\phi=2$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
@@
Figure 5: (a) Spectral functions $A_{z}(\omega)$ (solid lines) and occupation functions $N_{z}(\omega)$ (shaded areas and dashdotted lines) for the different layers at $\Omega=8$, $\Delta V=5$ and $\lambda_{\textrm{eff}}=0.752$. (b) Zoom of electron DOS and occupation function for the layer $z=1$. Vertical dashed and dashdotted lines in (a) highlight the separation between the center of the bands of the layers $z=1$ and $z=4$, without (red line) and with e-ph interaction (coral line). Vertical black lines in (b) highlight the difference in the peak positions of the bands without and with e-ph interaction. Default parameters are specified in Table 1. (Here $\phi=2$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
IV.1.1 The Hartree shift

### v2

@@ 3.2 Observables
with $t_{\mathrm{rel}}=t-t^{\prime}$ and $t_{\mathrm{av}}=(t+t^{\prime})/2$ the relative and average times, respectively. The mean $j$  [^3] As discussed in Ref. [31], in a steady-state situation $j_{\tilde{z},\tilde{z}+1}$ should have the same value for every $\tilde{z}$. Nevertheless, due to the limited accuracy of our AMEA impurity solver [40], deviations occur. For this reason, we consider the mean $j$ of the photocurrent (averaged over the bonds), together with its standard deviation $\sigma_{j}$. of the steady-state photocurrent flowing from the left to the right lead [40, 36, 31] and its corresponding standard deviation $\sigma_{j}$ are
$j$ $\equiv\frac{1}{\mathrm{L}+1}\sum_{\tilde{z}=0}^{\mathrm{L}}j_{\tilde{z},\tilde{z}+1},$
$\sigma_{j}$ $\equiv\sqrt{\sum_{\tilde{z}=0}^{\mathrm{L}}\frac{1}{\mathrm{L}+1}(j_{\tilde{z},\tilde{z}+1}-j)^{2}},$ (15)
where $j_{\tilde{z},\tilde{z}+1}$ is the photocurrent between any of the $(\mathrm{L}+1)$ bonds of the entire Mott photovoltaic device [^4] The bonds take into account also the one from the left lead ($\tilde{z}=0$) to the first layer ($\tilde{z}=1$) and from the last layer ($\tilde{z}=\mathrm{L}$) to the right lead ($\tilde{z}=\mathrm{L}+1$)., given by:
$j_{\tilde{z},\tilde{z}+1}=t_{\tilde{z},\tilde{z}+1}\int_{-\infty}^{+\infty}\frac{\mathop{}\!\mathrm{d}\omega}{2\pi}\int\mathop{}\!\mathrm{d}\epsilon\int\mathop{}\!\mathrm{d}\overline{\epsilon}\rho\mathop{\mathrm{Re}}\nolimits(J_{00,\tilde{z},\tilde{z}+1}),$ (16)
where $\rho$  [^5] In Eq. 16 we omit the ($\omega,\epsilon,\varepsilon$) dependence to facilitate the reading. is the JDOS from Sec. 2 and the index $\tilde{z}$ runs over the leads’ surface layers and the correlated region, i.e. $\tilde{z}\in\left\{0,1,\dots,\mathrm{L}\right\}$ with $\tilde{z}=0$ indicating the left and $\tilde{z}+1=\mathrm{L}+1$ the right lead surface, the GF of which are given in Eq. 3.1.
We characterize the spectral properties of this system via the local electron density of states (DOS) $A_{z}(\omega)$ and occupation function $N_{z}(\omega)$  [^6] The occupation function $N_{z}(\omega)$ tells us how many electronic states per spin are occupied with energy $\omega$ in the layer $z$ and is proportional to $G^{<}_{\mathrm{loc},00,zz}$., defined as
$A_{z}(\omega)=-\frac{1}{\pi}\mathop{\mathrm{Im}}\nolimits[G_{\mathrm{loc},00,zz}^{\mathrm{R}}(\omega)],$ (17)
@@
$N_{z}(\omega)=\frac{1}{4\pi}\left\{\mathop{\mathrm{Im}}\nolimits\left[G^{\mathrm{K}}_{\mathrm{loc},00,zz}(\omega)\right]-2\mathop{\mathrm{Im}}\nolimits\left[G^{\mathrm{R}}_{\mathrm{loc},00,zz}(\omega)\right]\right\},$ (18)
in which $G^{\mathrm{R}}_{\mathrm{loc},00,zz}(\omega)$ and $G^{\mathrm{K}}_{\mathrm{loc},00,zz}(\omega)$ represent respectively the time-averaged retarded and Keldysh local GF, introduced in A. Another important quantity we consider is the layer-dependent double occupation
$N_{\mathrm{D},z}=\langle\hat{n}_{z,\uparrow}\hat{n}_{z,\downarrow}\rangle,$ (19)
@@
$\Delta N_{\mathrm{D},z}$ $=\langle\left(\hat{n}_{z,\uparrow}-\langle\hat{n}_{z,\uparrow}\rangle\right)\left(\hat{n}_{z,\downarrow}-\langle\hat{n}_{z,\downarrow}\rangle\right)\rangle$ (20)
$=N_{\mathrm{D},z}-n_{z}^{2}$
and a mean-field contribution $n_{z}^{2}$, with $n_{z}=\langle\hat{n}_{z,\uparrow}\rangle=\langle\hat{n}_{z,\downarrow}\rangle$ denoting the number of particles per spin per layer  [^7] In this work, we consider the central correlated system in a paramagnetic phase. As mentioned in Sec. 1, antiferromagnetic correlations are important for the spreading of photoexcited carriers [38] and should be included in the model for a possible enhancement of the efficiency of a Mott photovoltaic system.. $n_{z}$ enters also into the total exceeding charge in the layer $z$
@@
A quantity related to the efficiency of the Mott photovoltaic device, which is discussed in Sec. 4, is the time-averaged absorbed power $\overline{P}_{\mathrm{abs}}$
$\overline{P}_{\mathrm{abs}}=-E_{0}\int\mathop{}\!\mathrm{d}\epsilon\int\mathop{}\!\mathrm{d}\overline{\epsilon}\rho(\epsilon,\overline{\epsilon})\sum_{l=-\infty}^{\infty}\Big{(}J_{1-l}(\mathscr{A})$
$+J_{-1-l}(\mathscr{A})\Big{)}\Big{[}\epsilon(1-\delta_{l,0})\delta_{l,2k}+\mathrm{i}\overline{\epsilon}\delta_{l,2k+1}\Big{]}$
$\sum_{z=1}^{\mathrm{L}}\int_{-\infty}^{+\infty}\frac{\mathop{}\!\mathrm{d}\omega}{2\pi}G_{l,z}^{<}(\omega;\epsilon,\overline{\epsilon})$ (22)
in which $k\in\mathbb{Z}$. We use in 3.2 the so-called Wigner representation, introduced in D, for the lesser GF (see Sec. 3.1).
4 Results
We consider the setup shown in Fig. 1, with the local DOS landscape in Fig. 2, consisting of $\mathrm{L}=4$ correlated layers, a situation which has already been studied in Refs. [43, 24, 31]. The default values for the main parameters employed in this manuscript are given in Table 1  [^8] The values of the parameters in Table 1 are such that $\alpha\equiv t^{\ast}_{\parallel}E_{0}/2\Omega^{2}<0.5$, which justifies FDSA [25, 30, 31]., if not stated otherwise.
$U$ $E_{0}$ $\gamma_{\mathrm{l}/\mathrm{r}}$ $v$ $t_{\mathrm{l}/\mathrm{r}}$ $t_{\perp}$ $\Phi$ $1/\beta$ $\mathrm{L}$
@@ 4 Results
Table 1: Default values of the main parameters used in this manuscript: for simplicity we defined $v\equiv v_{\mathrm{l}/\mathrm{r}}$. As a reminder, the renormalized hopping in the correlated region is $t^{\ast}_{\parallel}=2$.
@@
We choose the leads’ onsite energies $\varepsilon_{\mathrm{l}/\mathrm{r}}$ such that the left(right) lead DOS overlaps with the lower(upper) Hubbard band of the leftmost(righmost) layer with $\varepsilon_{\mathrm{l}}=\varepsilon_{1}$ and $\varepsilon_{\mathrm{r}}=-\varepsilon_{\mathrm{l}}$, see the scheme in Fig. 2.
@@
Figure 2: Schematic representation of impact ionization processes occurring in the system under analysis. Electrons are injected into the central insulating region (in green) by the left reservoir (in blue) and drained out of it by the right reservoir. $\Phi$ is the potential drop across the central region and is a linear function of the coordinate $z$ of the layers. The bias voltage $\Delta V=\mu_{r}-\mu_{l}$ applied to the metallic leads sets the occupations within them and produces a current from the right to the left of the heterostructure (orange arrows). For values of the driving frequency $\Omega\gtrsim 2\Delta_{\mathrm{g}}$, a photoexcited electron in the upper band can excite, in turn, a second electron across the gap from the lower band by transferring energy to it via Coulomb scattering. As outcome, two carriers per photon can now tunnel towards the right lead, resulting in a current from the left to the right (denoted by the red arrows), which flows against the potential energy established by the bias voltage.
@@
We vary the bias voltage $\Delta V=\mu_{r}-\mu_{l}$, so that for $\Delta V>0$ energy is harvested from the periodic driving when an electron current flows from the left to the right lead. We take, as convention, the current flowing from right to left as positive and the one from left to right as negative. The local DOS of the correlated layers has a gap $\Delta_{\mathrm{g}}\approx 3$, while the bandwidth of the leads $W_{\mathrm{b}}\approx 8.5$ [30, 31]. We quantify the effect of acoustic phonons on the electronic observables via the effective e-ph coupling strength $\lambda_{\mathrm{eff}}=2g^{2}/W_{\mathrm{b}}\omega_{ph}$, whose derivation is found in C  [^9] The values of $\lambda_{\mathrm{eff}}$ in Sec. 4 are obtained taking $W_{\mathrm{b}}=8.5$, $g\in\left\{0.3,0.4,0.5\right\}$ and $\omega_{ph}\in\left\{0.05,0.1\right\}$, in line with Ref. [30, 62]..
@@
In this work we are interested in assessing how II is influenced by the dissipation by acoustic phonons and how it affects the efficiency of the Mott photovoltaic system. This process, in which an electron excited to the upper Hubbard band (UHB) by a photon with energy $\Omega$ excites a second electron across the gap via electron-electron scattering, may occur if the bandwidth of the UHB, which in our setup is roughly equal to the leads’ DOS, is at least twice the size of the gap $\Delta_{\mathrm{g}}$, i.e. $W_{\mathrm{b}}\geq 2\Delta_{\mathrm{g}}$ [25, 30, 31].
4.1 Effects of electron-phonon interaction
@@
Figure 3: (a) Absolute value of the time-averaged steady state current $|j|$ with error bars corresponding to $\sigma_{j}$ and (b)-(c) fluctuation of double occupancy $\Delta N_{\mathrm{D},z}$ for $z=1$ and $z=2$, plotted as functions of the driving frequency $\Omega$, for different values of the bias voltage $\Delta V$ without and with e-ph interaction with coupling strength $\lambda_{\mathrm{eff}}=0.752$. Default parameters are specified in Table 1. Due to the PhI symmetry (see Ref. [31]), $\Delta N_{\mathrm{D},\mathrm{L}+1-z}=\Delta N_{\mathrm{D},z}$ so the curves for $z=3$ and $z=4$ in (b)-(c) are omitted. (Here $\phi=2$, $\Delta_{\mathrm{g}}\approx 3$ and $W_{\mathrm{b}}\approx 8.5$.)
@@
Fig. 3 shows that the photocurrent as well as $\Delta N_{\mathrm{D},z}$ for $\Delta V=2$ in presence of the e-ph interaction (green curves) display a slight enhancement for small to intermediate $\Omega$-values. This is due to the tiny redistribution of the spectral weight from the maxima peaks towards the edges of the Hubbard bands in Figs. 4, which reduces slightly the gap $\Delta_{g}$, with a larger occupation of the positive-energy states next to the bottom of the UHBs for all the layers. This effect is relatively small for a single layer but becomes appreciable when considering the whole heterostructure, and leads to a small increase of the photocurrent, as already observed in Ref. [30]. The magnitude of the peak in $j$ at $\Delta V=2$ remains instead comparable, because of the weaker DOS suppression around the Hubbard peaks [Fig. 4(b)] with respect to Ref. [30].
@@
We notice in Fig. 3(a) that the kink around $\Omega\approx 2\Delta_{g}$ signalizing II [31] is not affected by the e-ph interaction. In this regard, also the fluctuation of double occupancy $\Delta N_{\mathrm{D},z}$ [green curves in Fig. 3(b)-(c)] shows a change in slope for $\Omega\approx 2\Delta_{g}$, although less sharp due to the smearing by acoustic phonons.
@@
On the other hand, at $\Delta V=5$ the photocurrent experiences a reduction in magnitude because of the larger bias voltage which favours a current in the opposite direction  [^10] The current at $\Omega\gtrsim 17$ for $\Delta V=5$ becomes negative (not displayed in Fig. 3(a)), because for such $\Omega$ the carriers’ flow goes in the direction of the bias voltage.. The presence of acoustic phonons suppresses further $j$ because the Hartree shift, discussed in Sec. 4.1.1, moves the Hubbard bands away from each other, and the overlap between the layers’ DOS is reduced, as shown in Fig. 5(a).
@@
For small $\Omega$ the available states in the layers close to the right lead are mostly filled (see $N_{z}(\omega)$ for $z=3,4$ in Fig. 5(a)). Therefore, they suppress the photocurrent and contribute to a substantial increase of $\Delta N_{\mathrm{D},z}$ (coral curve in Fig. 3(b)-(c))  [^11] The residual value of $\Delta N_{\mathrm{D},z}$ at small and large $\Omega$ (red and coral curves in Fig. 3(b)-(c)) is due to the large separation between the Hubbard bands as a consequence of the Hartree shift..
@@
Figure 4: (a) Spectral functions $A_{z}(\omega)$ (solid lines) and occupation functions $N_{z}(\omega)$ (shaded areas and dashdotted lines) for the different layers at $\Omega=8$, $\Delta V=2$ and $\lambda_{\mathrm{eff}}=0.752$. (b) Zoom of electron DOS and occupation function for the layer $z=1$. Vertical dashed and dashdotted lines in (a) highlight the separation between the center of the bands of the layers $z=1$ and $z=4$, without (blue line) and with e-ph interaction (dark green line). Vertical black lines in (b) highlight the difference in the peak positions of the bands without and with e-ph interaction. Default parameters are specified in Table 1. (Here $\phi=2$, $\Delta_{\mathrm{g}}\approx 3$ and $W_{\mathrm{b}}\approx 8.5$.)
@@
Figure 5: (a) Spectral functions $A_{z}(\omega)$ (solid lines) and occupation functions $N_{z}(\omega)$ (shaded areas and dashdotted lines) for the different layers at $\Omega=8$, $\Delta V=5$ and $\lambda_{\mathrm{eff}}=0.752$. (b) Zoom of electron DOS and occupation function for the layer $z=1$. Vertical dashed and dashdotted lines in (a) highlight the separation between the center of the bands of the layers $z=1$ and $z=4$, without (red line) and with e-ph interaction (coral line). Vertical black lines in (b) highlight the difference in the peak positions of the bands without and with e-ph interaction. Default parameters are specified in Table 1. (Here $\phi=2$, $\Delta_{\mathrm{g}}\approx 3$ and $W_{\mathrm{b}}\approx 8.5$.)
4.1.1 The Hartree shift